### v4

# Quantum accessible information and classical entropy inequalities

Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. We show that the recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed. Via the optimality criterion, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
1 Introduction
The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, see e.g. [6], [16], [9], [11], and important in particular for applications in quantum cryptography. The problem was solved in some special (notably, symmetric) cases but in general remains open. In section (2) of this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information. It turns out that the optimality criterion when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality, as explained in section 3.
In this light, the conjecture of an information-optimal measurement for an ensemble of equiprobable equiangular states – quantum pyramid, formulated and numerically substantiated by Englert and Řeháček in [2] – is reconsidered in section 4 of the present paper. In the case of acute pyramids the corresponding parametric family of entropy inequalities (31) is proposed in subsections 4.1-4.3. The inequalities suggest the new tight lower bounds for the Shannon entropy of a discrete probability distribution as compared e.g. to the known bound from [13], theorem 3.11.
In the subsection 4.4 the case of obtuse pyramids is considered to get further insight and to formulate the hypotheses concerning the corresponding entropy inequalities (72), (80). Our approach allows also to derive the nonlinear equations (60) (resp. (77)) serving to determine the critical values of the parameter $t$ of the optimal observable corresponding to transition from moderately to strongly acute (resp, obtuse) pyramids.
In the subsection 4.5 we derive the tight entropy inequality (83) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids. The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and tentatively is the optimal one.
The proposed entropy inequalities may be more convenient for numerical check than the direct optimization, however proving them analytically is a challenge. In section 5, the suggestions for proofs proposed by the second author as a response to the hypotheses (31), (72), (80) put forward by the first author, are presented. Via the optimality criterion of section 2, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
2 Accessible information: a criterion for optimality
Let $\mathcal{H}$ be a finite dimensional Hilbert space, $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ the convex set of quantum states (density operators on $\mathcal{H}$). Ensemble of quantum states is a collection $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ where $\rho_{j}\in\mathfrak{S,}$ and $\pi=\left\{\pi_{j}\right\}$ is a probability distribution. Let $\mathcal{M}=\left\{M_{k}\right\}_{k=1,\dots,n}$ be an observable i.e. a collection of operators on $\mathcal{H}$ such that $M_{k}\geq 0,\,\sum_{k=1}^{n}M_{k}=I$ (unit operator). The joint distribution of “input” $j$ and “output” $k$ is then
$p_{jk}=\pi_{j}\mathrm{Tr\,}\,\rho_{j}M_{k}$
and the Shannon mutual information between $j$ and $k$ is (throughout the paper $\log$ denotes the binary logarithm)
$I(\mathcal{E},\mathcal{M})=\sum_{j,k}p_{jk}\log\frac{p_{jk}}{\pi_{j}p_{k}},$
with
$p_{k}=\sum_{j}p_{jk}=\mathrm{Tr\,}\,\overline{\rho}M_{k},$
where $\overline{\rho}=\sum_{j}\pi_{j}\rho_{j}$ is the average state of the ensemble $\mathcal{E}.$
Accessible information of the ensemble is defined as
$A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$ (1)
It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$ (if the matrices of $\rho_{j}$ in some basis are real then the number of components is reduced to $\frac{d(d+1)}{2}$). In particular, the supremum can be replaced with the maximum.
From now on we assume that the average state $\overline{\rho}$ is nondegenerate. Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
$\sigma_{k}=p_{k}^{-1}\overline{\rho}^{1/2}M_{k}\overline{\rho}^{1/2}=\left|\phi_{k}\rangle\langle\phi_{k}\right|,$
where
$|\phi_{k}\rangle=p_{k}^{-1/2}\overline{\rho}^{1/2}|\varphi_{k}\rangle,$
and observable $\mathcal{M}^{\prime}$ with the components
$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$ (2)
we have the same average state
$\overline{\sigma}=\sum_{k}p_{k}\sigma_{k}=\overline{\rho}$
and the new expression for accessible information:
$A(\mathcal{E})=\max_{\mathcal{E}^{\prime}:\overline{\sigma}=\overline{\rho}}I(\mathcal{E}^{\prime},\mathcal{M}^{\prime})$ (3)
with
$I(\mathcal{E}^{\prime},\mathcal{M}^{\prime})$ $=$ $I(\mathcal{E},\mathcal{M})=-\sum_{j}\pi_{j}\log\pi_{j}+\sum_{k,j}p_{k}p(j|k)\log p(j|k)$
$=$ $H(\pi)-\sum_{k}\mathrm{Tr\,}\,\left(p_{k}\sigma_{k}\right)K\left(\sigma_{k}\right),$
where $p(j|k)=\mathrm{Tr\,}\,\sigma_{k}M_{j}^{\prime},$
$K\left(\sigma\right)=-\sum_{j}M_{j}^{\prime}\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}.$ (4)
Therefore
$A(\mathcal{E})=H(\pi)-\min_{\mathcal{E}^{\prime}:\overline{\sigma}=\overline{\rho}}\sum_{k}\mathrm{Tr\,}\,\left(p_{k}\sigma_{k}\right)K\left(\sigma_{k}\right)$ (5)
Remark 1.
On the other hand, (3) is also the value of $\overline{\rho}$-constrained classical capacity $C(\mathcal{M}^{\prime},\overline{\rho})$ of the measurement channel corresponding to the observable $\mathcal{M}^{\prime}$ [8]. If $H$ is the Hamiltonian and $E$ is the energy limit then the energy-constrained capacity is
$C(\mathcal{M}^{\prime},E)=\max_{\overline{\rho}:\mathrm{Tr\,}\overline{\rho}H\leq E}C(\mathcal{M}^{\prime},\overline{\rho})$
and the full classical capacity is just $C(\mathcal{M}^{\prime})=\max_{\overline{\rho}}C(\mathcal{M}^{\prime},\overline{\rho}).$
Theorem 1.
The minimization problem
$\min_{\mathcal{E}^{\prime}:\overline{\sigma}=\overline{\rho}}\sum_{k}\mathrm{Tr\,}\,\left(p_{k}\sigma_{k}\right)K\left(\sigma_{k}\right)$ (6)
has the dual problem
$\max\left\{\mathrm{Tr\,}\overline{\rho}\Lambda:\Lambda^{\ast}=\Lambda,\,\Lambda\leq K(\sigma)\text{ for all }\sigma\in\mathfrak{S}\right\}.$ (7)
The following statements are equivalent:
$\Lambda_{0}$ is the solution of the problem (7); ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ with $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|$ is the solution of the problem (6) (equivalently, observable $M_{k}^{0}=\left|\varphi_{k}^{o}\rangle\langle\varphi_{k}^{0}\right|,\,k=1,\dots,n,$ with
$|\varphi_{k}^{0}\rangle=\left(p_{k}^{0}\right)^{1/2}\overline{\rho}^{-1/2}|\phi_{k}^{0}\rangle,$
maximizes the accessible information (1));
a.
$\Lambda_{0}\leq K(\sigma)$ for all $\sigma\in\mathfrak{S};$
b.
$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]\,|\phi_{k}^{0}\rangle=0,\,k=1,\dots,n$.
a.
The entropy inequality
[…]
holds for all unit vectors $\psi\in\mathcal{H};$
b.
The unit vectors $|\phi_{k}^{0}\rangle$ turn (8) into equality:
[…]
The solution of the dual problem (7) is unique on the support of the operator $\overline{\rho}$ i.e. $\Lambda_{0}\overline{\rho}=\Lambda_{0}^{\prime}\overline{\rho}$ for any two solutions $\Lambda_{0},\Lambda_{0}^{\prime}.$ In particular, if $\overline{\rho}$ is non-degenerate, then the solution of the dual problem is unique.
According to (5), the value of the accessible information (5) is then
$A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$ (10)
Proof.
We introduce the POVM on $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ as
$\Pi(d\sigma)=\sum_{k}\left(p_{k}\sigma_{k}\right)\delta_{\sigma_{k}}(d\sigma),$
where $\delta_{\sigma_{k}}(d\sigma)$ is the $\delta$-measure concentrated in the point $\sigma_{k}\in\mathfrak{S.}$ then the problem (6) can be written as [^1] We refer to [7] for rigorous general definition of the traced integral and formulations of the optimality conditions.
$\mathrm{Tr}\int_{\mathfrak{S}}\,K(\sigma)\Pi(d\sigma)$ $\longrightarrow$ $\min$ (11)
[…]
$\Pi(\mathfrak{S})$ $=$ $\overline{\rho},$
which is similar to the Bayes estimation problem [6].This is a convex programming problem for which the dual problem is (7) and the optimality conditions are (ii.a) and
$\int_{B}\left[K(\sigma)-\Lambda_{0}\right]\Pi_{0}(d\sigma)=0,\quad B\subset\mathfrak{S}(\mathcal{H}),$
see [7] for detail. In terms of the optimal ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ the last condition can be rewritten as
$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]\,\sigma_{k}^{0}=0,\quad(p_{k}^{0}>0).$ (12)
Note that by summing over $k$ we obtain
$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\sigma_{k}^{0}\,=\Lambda_{0}\overline{\rho},$ (13)
where $\Lambda_{0}$ must be Hermitean operator. This implies that $\Lambda_{0}\overline{\rho}$ is the same for all solutions of the dual problem.
By taking into account that $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|,$ we can rewrite (12) as
$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]|\phi_{k}^{0}\rangle=0$ (14)
and (13) as
$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\Lambda_{0}\overline{\rho}.$ (15)
Taking into account (4), the condition (ii.a) becomes
$\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,\quad\psi\in\mathcal{H},$
for all $\sigma\in\mathfrak{S.}$ Without loss of generality we can assume that $\left\|\psi\right\|=1$ here. Then for fixed $\psi$ the minimal value of the lefthand side is
[…]
since the difference is just the relative entropy of two probability distributions.
Thus we arrive to the entropy inequality (8) as an equivalent of the condition (ii.a). Further, by taking into account that $K(\sigma_{k}^{0})-\Lambda_{0}\geq 0,$ the condition (ii.b) in the form (14) is the same as
$\langle\phi_{k}^{0}|K(\sigma_{k}^{0})|\phi_{k}^{0}\rangle=\langle\phi_{k}^{0}|\Lambda_{0}|\phi_{k}^{0}\rangle$
which via (4) means that the unit vectors $|\phi_{k}^{0}\rangle$, satisfying the constraint $\sum_{k}p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\overline{\rho},$ turn (8) into the equality (9). ∎
Especially important for us is the case of pure-states ensemble with $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ where $|\psi_{j}\rangle;j=1,\dots,d$, are linearly independent vectors. In this case the resolution of the identity (2) is
$M_{j}^{\prime}=|e_{j}\rangle\langle e_{j}|,$
where the vectors
$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$ (16)
by necessity form an orthonormal basis (depending on the ensemble $\mathcal{E}$). Then the entropy inequality (8) takes the form
$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (17)
where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1.$ (18)
The left-hand side is just the Shannon entropy of the p.d. $\left\{\left|z_{j}\right|^{2};j=1,\dots,d\right\}$, and such an inequality can be approached with the tools of classical analysis. We will need the following
Lemma 1.
Let $\lambda_{jk}$ be real numbers. If the inequality
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k}\leq 0$ (19)
holds for all real $z_{j}$ satisfying the condition (18), then it holds for all complex $z_{j}$ satisfying the same condition.
Proof.
Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (18) amounts to
$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1.\quad$ (20)
Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, and $\lambda_{jk}$ are real hence
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
$+$ $P\sum_{j=1}^{d}\lambda_{jk}x_{j}x_{k}+(1-P)\sum_{j=1}^{d}\lambda_{jk}y_{j}y_{k}$
$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\sum_{j=1}^{d}\lambda_{jk}x_{j}x_{k}\right]$
$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\sum_{j=1}^{d}\lambda_{jk}y_{j}y_{k}\right]\leq 0,$
because
$\sum_{j=1}^{m}\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}=\sum_{j=1}^{m}\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}=1,\quad$
and by assumption (19) holds for real $z_{j}$ satisfying (18). ∎
The optimality conditions can be useful in the following ways:
  1. 1.
    for a problem which has a hypothetical solution, the conditions can be used to verify its actual optimality. One inserts the conjectured $p_{k}^{0},\sigma_{k}^{0}$ into (12) to compute Hermitean $\Lambda_{0}\$. Then the main difficulty will be proof of the entropy inequality (8). In the continuous variable version, this was our strategy in the solution of the related Gaussian optimizers conjecture for the capacity of Gaussian measurement channels [8], with the new generalizations of the famous log-Sobolev inequality as the entropy inequality.
  2. 2.
    for a problem which has a known solution, we get an entropy inequality (8) as a consequence of the optimality. Also one can obtain an alternative proof of the optimality if one can find an independent proof of the entropy inequality. In the next section we illustrate this point with a simple example resulting in the yet nontrivial tight lower bounds for the binary Shannon entropy.
  3. 3.
    The uniqueness of the solution of the dual problem (assuming $\bar{\rho}$ non-degenerate) has the following useful consequence. Assume that the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ is symmetric in that $\pi_{j}\equiv 1/m$ and there is a group of *-automorphisms $\left\{\alpha_{g};g\in G\right\}$ of the algebra of operators on $\mathcal{H}$ acting transitively on the set of states $\left\{\rho_{j}\right\}_{j=1,\dots,m}$. Then $\Lambda_{0}$ is an invariant of this group: $\alpha_{g}^{\ast}\left[\Lambda_{0}\right]=\Lambda_{0};g\in G.$
3 Example: two pure states
The case of ensemble of two states was treated previously in [12], [10]. For simplicity of demonstration we take two pure equiprobable states. Let $|e_{0}\rangle,|e_{1}\rangle$ be an orthonormal basis in two-dimensional Hilbert space. Consider ensemble $\mathcal{E}=\left\{\pi_{\pm},\rho_{\pm}\right\},$ where $\pi_{\pm}=\frac{1}{2},\rho_{\pm}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|$ with $|\psi_{\pm}\rangle=\cos\alpha/2|e_{0}\rangle\pm\sin\alpha/2|e_{1}\rangle,\,0\leq\alpha\leq\pi.$ As argued in [12] the accesible information
$A(\mathcal{E})=1-h\left(\frac{1+\sin\alpha}{2}\right)$ (21)
is attained by the optimal observable $\mathcal{M=}\left\{M_{+},M_{-}\right\}$ where $M_{\pm}=\left|e_{\pm}\rangle\langle e_{\pm}\right|$ and $|e_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|e_{0}\rangle\pm|e_{1}\rangle\right)$ is an orthonormal basis [^2] Strictly speaking, the paper [12] contains the check of local optimality.. Here
$h(t)=-t\log t-(1-t)\log(1-t)$
is the classical binary entropy.
Let us apply theorem 1 giving a proof of the global optimality and, as a byproduct, a new entropy inequality. The average state of the ensemble $\mathcal{E}$ is
$\overline{\rho}=\frac{1}{2}\left|\psi_{+}\rangle\langle\psi_{+}\right|+\frac{1}{2}\left|\psi_{-}\rangle\langle\psi_{-}\right|=\cos^{2}\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin^{2}\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$
with
$\overline{\rho}^{1/2}=\cos\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$ (22)
so that
$\frac{1}{\sqrt{2}}\overline{\rho}^{-1/2}|\psi_{\pm}\rangle=|e_{\pm}\rangle$ (23)
and
$M_{\pm}^{\prime}=\frac{1}{2}\overline{\rho}^{-1/2}\rho_{\pm}\overline{\rho}^{-1/2}=\left|e_{\pm}\rangle\langle e_{\pm}\right|,$ (24)
which is also the conjectured optimal measurement. Therefore the conjectured optimal ensemble $\mathcal{E}_{0}^{\prime}$ coincides with $\mathcal{E}$ because
$\sigma_{\pm}^{0}=2\overline{\rho}^{1/2}\left|e_{\pm}\rangle\langle e_{\pm}\right|\overline{\rho}^{1/2}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|,\quad\pi_{\pm}^{0}=\frac{1}{2},$
 so that $|\phi_{\pm}^{0}\rangle=|\psi_{\pm}\rangle.$ Let us check the optimality conditions for the ensemble $\mathcal{E}^{0}=\mathcal{E}$.
Introducing the important parameter
$p=\frac{1+\sin\alpha}{2},\quad\frac{1}{2}\leq p\leq 1,$ (25)
we have
$\left\langle e_{\pm}|\psi_{\pm}\right\rangle=\sqrt{p},\quad\left\langle e_{\pm}|\psi_{\mp}\right\rangle=\sqrt{1-p}.$
Incerting $\sigma_{\pm}^{0}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|$ into
[…]
we obtain
$K(\sigma_{+}^{0})$ $=$ $-\left|e_{+}\rangle\langle e_{+}\right|\log p-\left|e_{-}\rangle\langle e_{-}\right|\log(1-p),$
$K(\sigma_{-}^{0})$ $=$ $-\left|e_{+}\rangle\langle e_{+}\right|\log(1-p)-\left|e_{-}\rangle\langle e_{-}\right|\log p.$
The ensemble $\mathcal{E}$ is symmetric with respect to the reflection relative to the axis $|e_{0}\rangle;$ according to the remark 3 at the end of previous section, $\Lambda_{0}$ should also obey this symmetry, hence
$\Lambda_{0}=\lambda_{0}(p)\left|e_{0}\rangle\langle e_{0}\right|-\lambda_{1}(p)\left|e_{1}\rangle\langle e_{1}\right|.$ (26)
Solving the conditions (ii.b)
$\left[K(\sigma_{\pm}^{0})-\Lambda_{0}\right]\,|\phi_{\pm}^{0}\rangle=0$
and taking into account that
$\cos\alpha/2=\frac{1}{\sqrt{2}}\left(\sqrt{p}+\sqrt{1-p}\right),\quad\sin\alpha/2=\frac{1}{\sqrt{2}}\left(\sqrt{p}-\sqrt{1-p}\right),$
results in
$\lambda_{0}(p)=-\frac{\sqrt{p}\log p+\sqrt{1-p}\log\left(1-p\right)}{\sqrt{p}+\sqrt{1-p}},$
$\lambda_{1}(p)=\frac{\sqrt{p}\log p-\sqrt{1-p}\log\left(1-p\right)}{\sqrt{p}-\sqrt{1-p}},$
provided $p>1/2$.
Then the conditions (ii.b) are fulfilled while (ii.a) reduces to (8) which via (24), (26) amounts to the following parametric family of tight lower bounds for the classical binary entropy  [^3] See the dynamical plot of the inequality at https://www.desmos.com/Calculator/11ikk1i8fb
Proposition 1.
For $p>1/2$ and $0\leq t\leq 1$ (in our case $t=|\left\langle e_{+}|\psi\right\rangle|^{2},1-t=|\left\langle e_{-}|\psi\right\rangle|^{2}$),
$h(t)\geq(\lambda_{0}(p)+\lambda_{1}(p))\sqrt{t(1-t)}-\frac{\lambda_{1}(p)-\lambda_{0}(p)}{2},$ (27)
where
$\frac{\lambda_{0}(p)+\lambda_{1}(p)}{2}=\sqrt{p\left(1-p\right)}\frac{\log p-\log\left(1-p\right)}{2p-1}\equiv\mu_{0}(p),$
$\frac{\lambda_{1}(p)-\lambda_{0}(p)}{2}=\frac{p\log p-\left(1-p\right)\log\left(1-p\right)}{2p-1}.$
The inequality (27) becomes equality for $t=p$ and t=1p.t=1-p.
In these relations $p>1/2;$ taking the limit $p\rightarrow 1/2$ the inequality (27) becomes
$h(t)\geq 2\log e\sqrt{t(1-t)}-\log\frac{e}{2}.$ (28)
The value $p=1/2$ corresponds to the trivial case $\rho_{+}=\rho_{-}$ with $A(\mathcal{E})=0$. It is remarkable, however, that in the limit $p\rightarrow 1/2$ we get the nontrivial entropy inequality (28). It is equivalent to an inequality which was a starting point in the original derivation of logarithmic Sobolev inequality by L. Gross [3].
Thus (27), (28) can be regarded as discrete “distant relatives” of the log-Sobolev inequality. In the paper [8] we elaborated continuous variable version of the optimality conditions and applied it in the bosonic Gaussian case; the corresponding entropy inequality turned out to be a generalization of the genuine log-Sobolev inequality.
Proof.
For convenience of differentiation we pass to natural logarithms. Denote by $g(t,p)$ the righthand side of (27), then
$\frac{d}{dt}h(t)$ $=$ $-\ln t+\ln(1-t),$ (29)
$\frac{d}{dt}g(t,p)$ $=$ $\mu_{0}(p)\frac{1-2t}{\sqrt{t\left(1-t\right)}}.$ (30)
Taking into account the symmetry of the functions with respect to the point $t=1/2,$ we can restrict to the interval $t\in[1/2,1].$ One can check that $h(p)=g(p,p)$ and $\frac{d}{dt}|_{t=p}h(t)=\frac{d}{dt}|_{t=p}g(t,p)$ which means that $h(t)$ is the envelop of the family $g(t,p);\,p\geq 1/2,$ with $t=p$ as the tangential points. By using (29), (30) we obtain the key relation
$\frac{d}{dt}\left[h(t)-g(t,p)\right]=\left[\mu_{0}(p)-\mu_{0}(t)\right]\frac{2t-1}{\sqrt{t\left(1-t\right)}}.$
The function
$\mu_{0}(t)=\sqrt{t\left(1-t\right)}\frac{\ln t-\ln\left(1-t\right)}{2t-1}$
is monotonously decreasing on $[1/2,1],$ which can be seen by showing its derivative is negative on $(1/2,1]$ (for this it is convenient to make a monotonic change of variable $u=\sqrt{t/\left(1-t\right)}$).
Thus h(t)g(t,p)=0h(t)-g(t,p)=0 for $t=p,$ and $\frac{d}{dt}\left[h(t)-g(t,p)\right]$ has the same sign as tp.t-p. Hence $h(t)-g(t,p)\geq 0.$
Theorem 1 then implies the global optimality of the observable $\mathcal{M=}\left\{M_{+},M_{-}\right\}.$ Also
$\mathrm{Tr\,}\overline{\rho}\Lambda_{0}=\lambda_{0}(p)\cos^{2}\alpha/2-\lambda_{1}(p)\sin^{2}\alpha/2=h(p)$
and (10) implies $A(\mathcal{E})=1-h(p)$ in accordance with (21).
4 Multidimensional case
4.1 The lower bound for the Shannon entropy: formulation and discussion
The multidimensional generalization of (27) to be derived in what follows is:
Theorem 2.
For integer $m\geq 2$ and $p>\frac{m-1}{m}$
$\sum_{j=1}^{m}t_{j}\log t_{j}\geq\mu_{0}(p)\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\mu_{1}(p),$ (31)
where $t_{j}\geq 0,\sum_{j=1}^{m}t_{j}=1$ and
$\mu_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\sqrt{p}+\left(m-1\right)\sqrt{\frac{1-p}{m-1}}\right)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)},$ (32)
$\mu_{1}(p)=\frac{\sqrt{p}\log p-\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)}.$ (33)
This lower bound is tight: the graphs of the functions of $t_{1},\dots,t_{m}$ in the left and in the right sides of (31) are tangential at the points obtained by permutations of $t_{1}=p,t_{k}=\frac{1-p}{m-1},k=2,\dots,m$.
The inequality (31) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection. The suggestion for a proof of (31) will be given in section 5.1. Note that in the case $m=2$ the inequality (31) reduces to (27).
For $m\geq 3$ the value $p=\frac{m-1}{m}$ is allowed giving the inequality
$-\sum_{j=1}^{m}t_{j}\log t_{j}$ $\geq$ $\frac{\log\left(m-1\right)}{m-2}\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\left[\frac{m\log\left(m-1\right)}{m-2}-\log m\right]\quad$ (34)
$=$ $\log m-\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right],$ (35)
where
$B(P;P_{U})=\sum_{j=1}^{m}\sqrt{t_{j}/m}$
is the Bhattacharyya coefficient between the probability distribution $P=\left(t_{1},\dots,t_{m}\right)$ and the uniform distribution $P_{U}=\left(1/m,\dots,1/m\right).$
The inequality (35) then gives a lower bound for the Shannon entropy in terms of a “distance” between the distribution $P$ and the uniform distribution $P_{U}$, related to the Bhattacharyya coefficient. It can be compared with the lower bound of theorem 3.11 in [13]:
$\sum_{j=1}^{m}t_{j}\log t_{j}\geq\log m-\frac{m\log m}{2\left(m-1\right)}D(P;P_{U}),$ (36)
where $D(P;P_{U})=\sum_{j=1}^{m}\left|t_{j}-1/m\right|$ is the total variation distance. There is well-known relation for any two distributions $P,P^{\prime}$:
$2\left(1-B(P;P^{\prime})\right)\leq D(P;P^{\prime})\leq 2\sqrt{1-B(P;P^{\prime})^{2}},$
which allows to obtain from (35) a degraded version of (36) and vice versa. For $P$ close to $P_{U},$ the bound (35) is substantially better than (36) while close to degenerate distributions with $B(P;P_{U})\approx\sqrt{1/m}$ the right-hand side of (35) becomes negative which never happens for (36).
Similarly to (36), the inequality (35) (as well as (31)) can be rewritten as a tight upper bound for the classical relative entropy (the Kullback-Leibler divergence)
$H(P||P_{U})\leq\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right].$
4.2 The conjecture: “quantum pyramids”
Let $|e_{j}\rangle,\,j=1,\dots,m,$ be an orthonormal basis in $m$-dimensional Hilbert space.
$|e_{0}\rangle=\frac{1}{\sqrt{m}}\sum_{j=1}^{m}|e_{j}\rangle,\quad\langle e_{0}|e_{j}\rangle\equiv\frac{1}{\sqrt{m}}.$
Following Englert and Řeháček [2] we consider the system of equiangular vectors – the quantum pyramid:
$|\psi_{j}\rangle=\sqrt{mr_{1}}|e_{j}\rangle+\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)|e_{0}\rangle,\quad j=1,\dots,m,$ (37)
$r_{0},\,r_{1}\geq 0,\quad r_{0}+(m-1)r_{1}=1.$ (38)
Alternative parametrization was given in [4], section VI.2, where this configuration was related to “orthogonal signals” in communication theory:
$|\psi_{j}\rangle=a|e_{j}\rangle+b\sum_{k\neq j}|e_{k}\rangle,\quad j=1,\dots,m,$ (39)
$a\geq 0,\quad a^{2}+(m-1)b^{2}=1.$ (40)
The relation between the two parametrizations is
$a=\sqrt{mr_{1}}+\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},\quad b=\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},$ (41)
$1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}.$ (42)
The cosine between the edges is $\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}\,,\,j\neq k,$ thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute, $r_{0}=r_{1}\,(b=0)$ – to the orthogonal, and $r_{0}<r_{1}\,(b<0)$ – to the obtuse pyramids. The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle,$ and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat pyramid.
We will consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
$\bar{\rho}\equiv\frac{1}{m}\sum_{j=1}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (43)
Remark 2.
Operators of the form $c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|$ constitute Abelian algebra of operators invariant under the group of all *-automorphisms generated by the orthogonal transformations leaving the quantum pyramid invariant.
One can then easily check that
$\bar{\rho}^{1/2}=\sqrt{r_{1}}I+\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)|e_{0}\rangle\langle e_{0}|$ (44)
and
$|\psi_{j}\rangle=\sqrt{m}\bar{\rho}^{1/2}|e_{j}\rangle.$ (45)
In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2]. In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form $\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
$|\tilde{e}_{k}\rangle=|e_{k}\rangle+\frac{t-1}{\sqrt{m}}|e_{0}\rangle;\quad k=1,\dots,m,$ (46)
$|\tilde{e}_{0}\rangle=\sqrt{1-t^{2}}|e_{0}\rangle.$
The optimal value of the parameter is
$t_{a}(m)=\left\{\begin{array}[]{cc}1,&r_{0}\leq\frac{4(m-1)}{m^{2}}\\ \frac{2(m-1)}{m-2}\sqrt{\frac{r_{1}}{r_{0}}},&r_{0}>\frac{4(m-1)}{m^{2}}\end{array}\right$., (47)
giving the accessible information (eq. (36) in [2])
$A(\mathcal{E})=\left\{\begin{array}[]{cc}I^{(SRM)},&r_{0}\leq\frac{4(m-1)}{m^{2}}\\ \frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),&r_{0}>\frac{4(m-1)}{m^{2}}\end{array}\right$., (48)
where
$I^{(SRM)}$ $=$ $\frac{1}{m}\left(\sqrt{r_{0}}+\left(m-1\right)\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}+\left(m-1\right)\sqrt{r_{1}}\right)^{2}$ (49)
$+$ $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
$=$ $\log m+a^{2}\log a^{2}+\left(m-1\right)b^{2}\log b^{2}.$ (50)
Note that in the case $t_{a}(m)=1$ the conjectured information-optimal observable $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,\dots,m,$ is optimal for the Bayes discrimination problem for the ensemble $\mathcal{E}$ [4], sec. VI.2. It is called square-root measurement (SRM) since the system $\left\{|{e}_{k}\rangle\right\}$ is obtained from $\left\{|{\psi}_{k}\rangle\right\}$ by applying inverse square root of its Gram matrix, cf. (45).
Remark 3.
An upper bound and a benchmark for $A(\mathcal{E})$ is the classical capacity of the c-q channel $j\rightarrow|\psi_{j}\rangle\langle\psi_{j}|$:
$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$ (51)
The last equality follows from the symmetry of the quantum pyramid and concavity of the von Neumann entropy. According to (43) the density operator $\bar{\rho}$ has the eigenvalues $r_{0},r_{1}$, the last of multiplicity m1m-1, hence
$C$ $=$ $-r_{0}\log r_{0}-(m-1)r_{1}\log r_{1}$
$=$ $-\left(1-\frac{m-1}{m}(a-b)^{2}\right)\log\left(1-\frac{m-1}{m}(a-b)^{2}\right)$
- $\frac{m-1}{m}(a-b)^{2}\log\frac{(a-b)^{2}}{m},$
where in the second equality we used (42).
Unlike the case of accessible information, this expression is the same for all values of the parameters, for acute, obtuse and flat pyramids. The ratio $C/A(\mathcal{E})$ gives the gain due to the measurement entanglement at the channel’s output. In particular, it tends to the infinity as the pyramid collapses to ort (weak signal) and equals to 1 for rectangular pyramid (ideal channel).
4.3 Checking the optimality conditions: the case of acute pyramids
Let us apply our approach of section 2. By using (45) the dual observable $\mathcal{M}^{\prime}$ has the components
$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\frac{1}{m}|\psi_{j}\rangle\langle\psi_{j}|\right)\overline{\rho}^{-1/2}=|e_{j}\rangle\langle e_{j}|,$ (52)
hence
[…]
From (46), (44) we obtain the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
$\sigma_{k}=p_{k}^{-1}|\tilde{\phi}_{k}\rangle\langle\tilde{\phi}_{k}|\,,\quad p_{k}=\langle\tilde{\phi}_{k}|\tilde{\phi}_{k}\rangle$
where
$|\tilde{\phi}_{k}\rangle$ $=$ $\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle;$
$p_{k}$ $=$ $r_{1}+\frac{r_{0}t^{2}-r_{1}}{m};\quad k=1,\dots,m;$
$|\tilde{\phi}_{0}\rangle=\overline{\rho}^{1/2}|\tilde{e}_{0}\rangle=\sqrt{r_{0}\left(1-t^{2}\right)}|e_{0}\rangle;\quad p_{0}=r_{0}\left(1-t^{2}\right).$
Substitution gives
$K\left(\sigma_{k}\right)=-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log p_{k}^{-1}\mathrm{\,}\,\left|\langle e_{j}|\tilde{\phi}_{k}\rangle\right|^{2},$
hence
$K\left(\sigma_{k}\right)=\left\{\begin{array}[]{cc}-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log p_{k}^{-1}\mathrm{\,}\,\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2},&k\geq 1\\ -\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log\frac{1}{m}=\left(\log m\right)I,&k=0\end{array}\right.$.
Having in mind checking the optimality condition (ii.b) we compute
$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$ $=$ $-\sum_{j}|e_{j}\rangle\left(\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$ (53)
$=$ $A(t)|e_{k}\rangle+B(t)\sqrt{m}|e_{0}\rangle;\quad k=1,\dots,m,$
where
$A(t)$ $=$ $-\left(\sqrt{r_{1}}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}-B(t),$ (54)
$B(t)$ $=$ $-\left(\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$ (55)
and
$K\left(\sigma_{0}\right)|\tilde{\phi}_{0}\rangle=\sqrt{r_{0}\left(1-t^{2}\right)}\log m\,|e_{0}\rangle.$
Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 we look for $\Lambda_{0}$ in the form
$\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|.$ (56)
Then
$\Lambda_{0}|\tilde{\phi}_{k}\rangle$ $=$ $c_{1}\sqrt{r_{1}}|e_{k}\rangle+\frac{\left(c_{0}+c_{1}\right)\sqrt{r_{0}}t-c_{1}\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle;\,k\geq 1,$
$\Lambda_{0}|\tilde{\phi}_{0}\rangle$ $=$ $\left(c_{0}+c_{1}\right)\sqrt{r_{0}\left(1-t^{2}\right)}|e_{0}\rangle.$
Comparing coefficients in the condition (ii.b): $\Lambda_{0}|\tilde{\phi}_{k}\rangle=K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$, we obtain the system of equations
$c_{1}\sqrt{r_{1}}$ $=$ $A(t),$
$\left(c_{0}+c_{1}\right)\sqrt{r_{0}}t-c_{1}\sqrt{r_{1}}$ $=$ $mB(t),$ (57)
$\left(c_{0}+c_{1}\right)\sqrt{1-t^{2}}$ $=$ $\log m\sqrt{1-t^{2}}.$
Two cases are possible: $t=1$ and $t<1.$ In the first case the third equation disappears while the first two turn into
$c_{1}\sqrt{r_{1}}$ $=$ $\frac{b\log b^{2}-a\log a^{2}}{\sqrt{m}},$
$\left(c_{0}+c_{1}\right)\sqrt{r_{0}}-c_{1}\sqrt{r_{1}}$ $=$ $-\sqrt{m}b\log b^{2},$
where we have used (54), (55) and (41). Taking into account the relations (41), we have $\sqrt{mr_{1}}=a-b,$ and the solution is
$c_{1}$ $=$ $\frac{b\log b^{2}-a\log a^{2}}{a-b},$
$c_{0}$ $=$ $m\frac{ab\left(\log a^{2}-\log b^{2}\right)}{\left(a-b\right)\left[a+\left(m-1\right)b\right]}.$ (58)
In the case $t<1$ the third equation in the system (57) reduces to $c_{0}+c_{1}=\log m.$ Eliminating $c_{0},c_{1}$ one obtains the equation for $t:$
$A(t)+mB(t)=\left(\log m\right)\sqrt{r_{0}}t.$ (59)
Substituting the expressions (54), (55) for $A(t),B(t)$ and introducing the parameter
$\tau=\sqrt{\frac{r_{0}}{r_{1}}}t,$
we can reduce (59) to the form
$\tau\log\left(m-1+\tau^{2}\right)=\frac{m-1+\tau}{m}\log\left(m-1+\tau\right)^{2}+\frac{m-1}{m}\left(\tau-1\right)\log\left(\tau-1\right)^{2}.$ (60)
Denoting $s=\tau^{-1}$ equation (60) can be transformed into
$f(s)$ $\equiv$ $\ln\left[1+\left(m-1\right)s^{2}\right]$
- $2\frac{1+\left(m-1\right)s}{m}\ln\left[1+\left(m-1\right)s\right]-2\frac{m-1}{m}\left(1-s\right)\ln\left|1-s\right|=0.$
At the point $s=0$ (corresponding to $\tau\rightarrow\infty$) this function and its first and second derivatives vanish, while $f^{\prime\prime\prime}(s)=2\left(m-1\right)\left(m-2\right)>0$ hence $f$ is increasing and positive for small $s$. But $f(1)=-\ln m<0,$ hence there is a (unique) root in the interval $(0,1],$ namely $s=\frac{m-2}{2\left(m-1\right)}.$ There are no roots for $s>1$ because the function is decreasing.
Thus the value $\tau_{a}(m)=\frac{2(m-1)}{m-2}$ suggested by (47) is the unique positive solution of (59). Substituting this value in the first equation of (57) and using the third equation we obtain
$c_{1}=\log m-\frac{m\log\left(m-1\right)}{m-2},\quad c_{0}=\frac{m\log\left(m-1\right)}{m-2},$ (61)
the values which do not depend on the parameters of the pyramid $r_{0},r_{1}.$ The criterion for the applicability of this second case is
$t_{a}(m)\equiv\sqrt{\frac{r_{1}}{r_{0}}}\tau_{a}(m)=\sqrt{\frac{r_{1}}{r_{0}}}\frac{2(m-1)}{m-2}<1$
which, via (38), is equivalent to $r_{0}>\frac{4(m-1)}{m^{2}}$, and, via (41), to $a^{2}<\frac{m-1}{m}.$
Correspondingly, the condition for applicability of the solution (58) is $a^{2}\geq\frac{m-1}{m}.$ By introducing the parameter $p$ such that
$p=a^{2},\quad\frac{1-p}{m-1}=b^{2},$ (62)
we have $b=\sqrt{\frac{1-p}{m-1}},$ leading to
$\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|=-\mu_{1}(p)I+\mu_{0}(p)m|e_{0}\rangle\langle e_{0}|$
with $\mu_{1}(p),\mu_{0}(p)$ given by (33), (32) under the condition that $p\geq\frac{m-1}{m}$. Introducing complex numbers $z_{j}=\langle e_{j}|\psi\rangle,$ so that $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,$ the optimality condition (ii.a) amounts to the inequality
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (63)
Since $\mu_{0}(p)\geq 0,$ this is equivalent to (31) where $t_{j}=\left|z_{j}\right|^{2}$ .
Remarkably, inserting the limiting value $p=\frac{m-1}{m}$ we obtain for $m>2$
$\mu_{1}=\frac{m\log\left(m-1\right)}{m-2}-\log m,\quad\mu_{0}=\frac{\log\left(m-1\right)}{m-2},$
which correspond to the values $c_{0},c_{1},$ given by (61) and obtained by solving (57) in the case $t_{a}(m)<1.$
Finally, from (56) and (43) $\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=c_{1}+c_{0}r_{0}$ so that
$A(\mathcal{E})=\log m-c_{1}-c_{0}r_{0}.$ (64)
Thus we have the following corollary to the theorem 2:
Corollary 1.
For ensemble of acute pyramids the information-optimal observable is given by the equations (46), (47) as conjectured in [2]. The accessible information is given by the expressions
$A(\mathcal{E})=\left\{\begin{array}[]{cc}\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1},&p\in\left[\frac{m-1}{m},1\right]\\ \frac{m-1}{m-2}\log(m-1)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)^{2},&p\in\left[\frac{1}{m},\frac{m-1}{m}\right]\end{array}\right$. (65)
which agree with (48).
Proof.
For accessible information we use the relation (64). A routine calculation using (58), (41), (62) gives (50) which is the first line in (65) for moderately acute pyramids. For strongly acute pyramid, using (61), one gets
$A(\mathcal{E})=\frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),$
which is the second line, taking into account (42). ∎
4.4 The case of obtuse pyramids
When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different. To describe it let us first consider the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf. [5]. Then the optimal observable is $\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [14]. In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
$\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (66)
which is saturated for $\alpha=\pi/2+\pi k$.
Most interesting is the case of “nearly flat” trine when $0<r_{0}<\bar{r}_{0}\approx 0,0614.$ Then, as shown by P. Shor in [15], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (37) with $r_{0}<r_{1}$ $(b<0)$ . The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
$M_{k}$ $=$ $t^{-2}|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|;\quad k=1,\dots,m;$ (67)
$M_{rs}$ $=$ $\frac{2}{m}\left(1-t^{-2}\right)|rs\rangle\langle rs|;\quad 1\leq r<s\leq m,$
where $t\geq 1,$ the vectors $|\tilde{e}_{k}\rangle$ are given by (46), and
$|rs\rangle=\frac{1}{\sqrt{2mr_{1}}}\left(|\psi_{r}\rangle-|\psi_{s}\rangle\right)=\frac{1}{\sqrt{2}}\left(|e_{r}\rangle-|e_{s}\rangle\right).$
Therefore the dual ensemble $\mathcal{E}^{\prime}$ consists of the states (we use the same notation as for similar objects in the acute case):
$\sigma_{k}=p_{k}^{-1}|\tilde{\phi}_{k}\rangle\langle\tilde{\phi}_{k}|\,,\quad p_{k}=\langle\tilde{\phi}_{k}|\tilde{\phi}_{k}\rangle,$
where
$|\tilde{\phi}_{k}\rangle$ $=$ $t^{-1}\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=t^{-1}\left(\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle\right),$
$p_{k}$ $=$ $t^{-2}\left(r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}\right);\quad k=1,\dots,m,$
and
$\sigma_{rs}=p_{rs}^{-1}|\tilde{\phi}_{rs}\rangle\langle\tilde{\phi}_{rs}|\,,\quad p_{rs}=\langle\tilde{\phi}_{rs}|\tilde{\phi}_{rs}\rangle,$
where
$|\tilde{\phi}_{rs}\rangle$ $=$ $\sqrt{\frac{2}{m}\left(1-t^{-2}\right)}\,\overline{\rho}^{1/2}|rs\rangle=\sqrt{\frac{r_{1}}{m}\left(1-t^{-2}\right)}\,\left(|e_{r}\rangle-|e_{s}\rangle\right);$
$p_{rs}$ $=$ $\frac{2r_{1}}{m}\left(1-t^{-2}\right).$
To check the optimality condition (ii.b) we proceed in parallel with the acute case, and compute
$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle=t^{-1}\left(A(t)|e_{k}\rangle+B(t)\sqrt{m}|e_{0}\rangle\right);\quad k=1,\dots,m,$
where $A(t),B(t)$ are given by (54), (55). Also
$K\left(\sigma_{rs}\right)|\tilde{\phi}_{rs}\rangle=-\sum_{j}|e_{j}\rangle\langle e_{j}|\tilde{\phi}_{rs}\rangle\,\log p_{rs}^{-1}\mathrm{\,}\,\left|\langle e_{j}|\tilde{\phi}_{rs}\rangle\right|^{2},$ (68)
with
$\langle e_{j}|\tilde{\phi}_{rs}\rangle$ $=$ $\sqrt{\frac{r_{1}}{m}\left(1-t^{-2}\right)}\left(\delta_{jr}-\delta_{js}\right),$
$\left|\langle e_{j}|\tilde{\phi}_{rs}\rangle\right|^{2}$ $=$ $\frac{r_{1}}{m}\left(1-t^{-2}\right)\left(\delta_{jr}+\delta_{js}\right).$
Substituting this into (68) results in
$K\left(\sigma_{rs}\right)|\tilde{\phi}_{rs}\rangle=|\tilde{\phi}_{rs}\rangle,\quad 1\leq r<s\leq m.$ (69)
Looking again for $\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|,$ and comparing coefficients in the condition (ii.b): $\Lambda_{0}|\tilde{\phi}_{k}\rangle=K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$, we obtain the first two equations of the system (57), namely
$c_{1}\sqrt{r_{1}}$ $=$ $A(t),$ (70)
$\left(c_{0}+c_{1}\right)\sqrt{r_{0}}t-c_{1}\sqrt{r_{1}}$ $=$ $mB(t),$
while $\Lambda_{0}|\tilde{\phi}_{rs}\rangle=K\left(\sigma_{rs}\right)|\tilde{\phi}_{rs}\rangle$ via (69) amounts to
$\left(1-t^{-2}\right)\left(c_{1}-1\right)=0,$ (71)
since $\langle e_{0}|\tilde{\phi}_{rs}\rangle\equiv 0.$
Again, two cases are possible: $t=1$ and $t>1.$ In the first case the third equation disappears while the first two lead to the same formulas (58) as in the acute case, however with $a=\sqrt{p\text{ }}$, but $b=-\sqrt{\frac{1-p}{m-1}}.$ Then, denoting $z_{j}=\langle e_{j}|\psi\rangle$, the condition (ii.a) brings us to the inequality:
For complex $z_{j}$ satisfying $\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ it holds
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (72)
where
$\tilde{\mu}_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\left(m-1\right)\sqrt{\frac{1-p}{m-1}}-\sqrt{p}\right)\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)},$ (73)
$\tilde{\mu}_{1}(p)=\frac{\sqrt{p}\log p+\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)}.$ (74)
The equality in (72) is attained for $z_{1}=\sqrt{p},\,z_{j}=-\sqrt{\frac{1-p}{m-1}},\,j=2,\dots,m,$ and all the permutations of these $z_{j}$.
Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$. However, as the lemma 1 shows, we can always restrict to the case of real $z_{j}$.
In (72) the parameter $p$ varies in the range $\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below. The accessible information in this case is given by the same expression (50) as in the case of moderately acute pyramids. For future use we note that
$\tilde{\mu}_{1}\left(\frac{m-1}{m}\right)=\log m-\frac{m-2}{m}\log(m-1).$ (75)
If $t>1,$ then (71) imply $c_{1}=1$ and equations (70) reduce to
$\sqrt{r_{1}}$ $=$ $A(t),$
$c_{0}$ $=$ $\frac{m}{\sqrt{r_{0}}t}B(t)+\frac{1}{t}\sqrt{\frac{r_{1}}{r_{0}}}-1.$ (76)
Here the first equation serves for determination of $t$ (compatible with the condition $t>1$). By using (54) it can be transformed to
$\log\frac{m\left(m-1+\tau^{2}\right)}{2}-\frac{m-1+\tau}{m}\log\left(m-1+\tau\right)^{2}-\frac{\left(1-\tau\right)}{m}\log\left(1-\tau\right)^{2}=0,$ (77)
where $\tau=\sqrt{\frac{r_{0}}{r_{1}}}t.$
Denoting $g(\tau)$ the function in equation (77), we have
$g(0)=\log\frac{m}{2}-\frac{m-2}{m}\log\left(m-1\right)\left\{\begin{array}[]{cc}>0,&m\leq 6;\\ <0,&m\geq 7,\end{array}\right.$.
The function is decreasing in the interval $[0,1]$ and $g(1)=-\log 2<0$, hence it has a root $\tau_{o}(m)$ in this interval if and only if $m\leq 6.$ For $\tau>1$ the function is negative and has no roots.
The condition $t>1$ amounts to $\tau_{o}(m)>\sqrt{\frac{r_{0}}{r_{1}}},$ which via (38), is equivalent to
$r_{0}<\left(1+\frac{m-1}{\tau_{o}(m)^{2}}\right)^{-1}\equiv r_{0}(m),$
and, via (41) and (62), to
$p\equiv a^{2}$ $<$ $\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}$ (78)
$=$ $\frac{(m-1+\tau_{o}(m))^{2}}{m(m-1+\tau_{o}(m)^{2})}\equiv p(m).$
One can check that equation (77) with $\tau=\tau_{o}(m)$ is equivalent to
$\tilde{\mu}_{1}(p(m))=1.$ (79)
The numerically found values [^4] By using equation (77) it is possible to express the function $p(m)$ via the Lambert $W$-function.:
$m$ 2 3 4 5 6 7
$\tau_{o}(m)$ 0 0,3616 0,25866 0,15195 0,04781 [-0,05286]
$mr_{0}(m)$ 0 0,1841 0,08726 0,02869 0,00274
$p(m)$ 0,5 0,8725 0,8656 0,85699 0,84896 0,8417
$\frac{m-1}{m}$ 0,5 0,6(6) 0,75 0,8 0.83(3) 0,8571
Here $mr_{0}(m)$ show a good agreement with the values found empirically for $m\leq 6$ in [2], eq. (39). In the case $m=7$, as well as for greater values of $m$, the solution of the equation (77) becomes negative and also $p(m)<\frac{m-1}{m}.$ This makes plausible the hypothesis that for $m\geq 7$ the optimal observable is SRM, and the corresponding entropy inequality is (72) for all $p\geq\frac{m-1}{m}.$
Then, for $m\leq 6$, under the condition (78) the equation (71) implies $c_{1}=1$ while from (76)
$c_{0}=\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}.$
The condition (ii.a) then amounts to the entropy inequality (72) with $\tilde{\mu}_{0}=c_{0}/m,\quad\tilde{\mu}_{1}=-1$ i.e.
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (80)
Due to lemma 1 one can restrict here to real $z_{j}$ , $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1.$ The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
In this case the accessible information, by using (64), is
$A(\mathcal{E})$ $=$ $\log\frac{m}{2}-r_{0}\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}$
$=$ $\log\dfrac{m}{2}-r_{0}m\tilde{\mu}_{0}(p(m)).$
The following theorem summarizes the content of this section.
Theorem 3.
If $m\geq 7$ then $p(m)<\frac{m-1}{m}$ and the entropy inequality (72) holds for $p\in\left[\frac{m-1}{m},1\right]$. If $m\leq 6$ then $p(m)>\frac{m-1}{m}$ and the inequality (72) holds for $p\in\left[p(m),1\right]$, yielding the inequality (80) for $p\in\left[\frac{m-1}{m},p(m)\right]$.
The suggestion for a proof of the inequalities is given in section 5.2.
Corollary 2.
In the case of obtuse pyramids the information-optimal observable is given by the relations (67) with $t=1$ for $m\geq 7$, and $t=\sqrt{\frac{r_{1}}{r_{0}}}\tau_{o}(m)$ for $m\leq 6$. The accessible information is given by the expression (50) if $p\in[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ otherwise by (4.4) if $m\leq 6$ and $p\in\left[\frac{m-1}{m},p(m)\right]$.
4.5 The case of flat pyramid
The considerations above formally do not include the case of flat pyramid with $r_{0}=0,r_{1}=1/(m-1).$ Then the average state of the ensemble $\mathcal{E}$
$\overline{\rho}=\frac{1}{m-1}\left(I-|0\rangle\langle 0|\right)$
is degenerate and we should use generalized inverse in the formula (52), resulting in
$M_{j}^{\prime}=|\tilde{\psi}_{j}\rangle\langle\tilde{\psi}_{j}|,\quad j=1,\dots,m,$
where $|\tilde{\psi}_{j}\rangle=|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle.$ Denoting $z_{j}=\langle e_{j}|\psi\rangle$ we have
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,\quad\sum_{j=1}^{m}z_{j}=0.$ (82)
Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$, hence the second equality follows.
Arguing as in the previous section, we suggest that the optimal observable is given by (67) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$. The corresponding tight entropy inequality is
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right$. (83)
for complex $z_{j}$ satisfying (82) (due to a simple modification of lemma 1 one can restrict here to real $z_{j}$ ). Indeed, for $m\leq 6$ we can use the argument leading to inequality (80) with $\sum_{j=1}^{m}z_{j}=0$ and $p=p(m)$. Then (79) implies the first line in the right-hand side of (83). On the other hand, for $m\geq 7$ we should use the inequality (72) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)\$gets the value (75) i.e. the second line in the righthand side of (83), which is greater than 1 for $m\leq 6$, but less for $m\geq 7$. We thus arrive at the following special case of theorem 3:
Proposition 2.
The tight entropy inequality (83) holds for real (or complex) $z_{j}$ satisfying the conditions (82).
In the case $m\leq 6$ the equality in (83) is attained for $z_{1}=-z_{2}=1/\sqrt{2},\,z_{j}=0$ for $j\geq 3;$ in the case $m\geq 7$ – for $z_{1}=\sqrt{\frac{m-1}{m}},\,z_{j}=-\sqrt{\frac{1}{\left(m-1\right)m}},\,j\geq 2,$ (and for all permutations of such $z_{j}).$
Theorem 1 then implies
Corollary 3.
For the ensemble of flat pyramid, information-optimal observable is given by (67) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$.
The corresponding accessible information is
$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\left\{\begin{array}[]{cc}\log\frac{m}{2},&m\leq 6;\\ \frac{m-2}{m}\log(m-1),&m\geq 7.\end{array}\right$.
For $m=3$ the inequality (83) reduces to (66). Indeed, as shown below, in (83) one can restrict to real $z_{j}$. Then $|\psi\rangle$ is a vector in the real plane $\langle e_{0}|\psi\rangle=0$ and
$z_{j}=\langle e_{j}|\psi\rangle=\langle\tilde{\psi}_{j}|\psi\rangle=\|\tilde{\psi}_{j}\|\,\cos\left(\alpha+\frac{2\pi j}{3}\right),\quad j=1,2,3,$
where $\||\tilde{\psi}_{j}\rangle\|=\||e_{j}\rangle-\sqrt{1/3}|e_{0}\rangle\|=\sqrt{2/3}$, hence (83) is the same as (66).
5 Suggestions for proofs
5.1 The case of theorem 2
Let $m\geq 2,\ \dfrac{m-1}{m}\leq p<1$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (32), (33).
For $p\in[(m-1)/m,1)$ and a probability distribution $(t_{1},...,t_{m})$, it holds
$F(t)\equiv H(t)-\Bigl(\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl)^{2}-\mu_{1}(p)\Bigl)\geq 0,$ (84)
where $H(t)=-\sum\limits_{k=1}^{m}t_{k}\log t_{k}$ is the Shannon entropy.
In the special case $p=1$, the inequality (84) is trivially satisfied by substituting $\mu_{0}(1)\equiv 0,\ \mu_{1}(1)=0$.
Proof.
The function $F(t)$ cannot reach a minimum on the boundary of the simplex of probability distributions, since the function
$u(s)=-\sum\limits_{i=1}^{m}(t_{*,i}+sa_{i})\log(t_{*,i}+sa_{i})-\mu_{0}(p)\Bigl(\sum\limits_{i=1}^{m}\sqrt{t_{*,i}+sa_{i}}\Bigl)+\mu_{1}(p)$
for a boundary point $t_{*}$ and a transversal direction $a$ has the derivative -$\infty$ at zero:
$u^{\prime}(0)=\lim\limits_{t\to t_{*}}\Bigg[-\sum\limits_{i\in I_{0}}a_{i}\log t_{i}-C\Bigl(\sum\limits_{i\in I_{0}}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl)+O(1)\Bigg],$
where $I_{0}=\{1\leq j\leq m;\ t_{*,j}=0\}$, $C>0$ is a constant. In this case, $a_{i}>0,\ i\in I_{0}$ (direction inside the simplex), hence $u^{\prime}(0)=$-$\infty$.∎
Thus it is sufficient to check the inequality (84) only at the critical points.
Lemma 2.
The critical points $t_{*}=(t_{*,1},...,t_{*,m})$ of the function $F(t)$ have the form
$t_{*}=\Bigl(q/k,...,q/k,(1-q)/(m-k),...,(1-q)/(m-k)\Bigl)$ (85)
up to permutations, for some integer $k$, $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$). One of the critical points is (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), at which the value of the function $F(t)$ is equal to 0.
Proof.
Consider the function $u(s)=F(t+sa)$, where $t=(t_{1},...,t_{m}),\ a=(a_{1},...,a_{m}),\ \sum\limits_{i}a_{i}=0$, for a range of $s$ such that $t+as$ defines a probability distribution. Then
$\dfrac{d}{ds}u(s)$ $=-\sum\limits_{i=1}^{m}a_{i}\log t_{i}-\mu_{0}(p)\Bigl(\sum\limits_{i=1}^{m}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl)\Bigl(\sum\limits_{i=1}^{m}\sqrt{t_{i}}\Bigl)$
$=\sum\limits_{i=2}^{m}a_{i}\Bigg[-\log t_{i}+\log t_{1}-\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl)\Bigl(\dfrac{1}{\sqrt{t_{i}}}-\dfrac{1}{\sqrt{t_{1}}}\Bigl)\Bigg].$
The condition $\dfrac{d}{ds}u(0)=0$ for a point $t_{*}=t$ and an arbitrary vector $a$ (that is, the point $t_{*}$ is critical) is satisfied if and only if for any $2\leq i\leq m$ it holds
$\log t_{*,1}-\log t_{*,i}=\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl)\Bigl(\dfrac{1}{\sqrt{t_{*,i}}}-\dfrac{1}{\sqrt{t_{*,1}}}\Bigl).$
Let us denote $C_{1}=\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl),C_{2}=\log t_{*,1}+\dfrac{C_{1}}{\sqrt{t_{*,1}}}$. The equation
$C_{2}-\log s=C_{1}/\sqrt{s},\quad s>0,$ (86)
has at most two roots, since the derivatives of the right and the left parts with respect to $s$ coincide only when $s=C_{1}^{2}/4$. But all the numbers $t_{*,1},...,t_{*,m}$ are solutions of the equation (86). Thus, among the values of $\{t_{*,i}\}$ there are no more than two different values, hence the lemma follows. ∎
At the critical points (85) the inequality $F(t_{*})\geq 0$ reduces to
$\mu_{0}(p,1/m)\Bigl(\sqrt{q(k/m)}+\sqrt{(1-q)(1-k/m)}\Bigl)^{2}-\mu_{1}(p,1-1/m)\quad$
$\leq$ $-q\log\dfrac{q}{k/m}-(1-q)\log\dfrac{1-q}{1-k/m},$
where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined as
$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big)}{\Bigl(\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$ (87)
$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (88)
for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
But this inequality is a special case of lemma 4 below, the proof of which is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$. We perform a qualitative analysis of the minimizer of this function. The global minimum (there can be no more than two local minima) appears on either the interval $[0,1/2]$ or $[1/2,1]$. Importantly, this segment contains no other critical points. Therefore, minimizing the function reduces to identifying the critical point within the appropriate interval and evaluating the function at that point. A detailed proof of this statement follows from
Lemma 3.
Let $A>0,\ B,\ C\in\mathbb{R}$ (the notation of the coefficients is used only within the framework of this lemma). Then the function
$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
[…]
Proof.
We have
$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
therefore local minima are on the interval $(0,1)$.
In the calculations we will pass to the natural logarithm.
$g^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl(\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl)-B=$
$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
Therefore, there are two inflection points located on different sides of $t=1/2$. It follows that there are three critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
The last statement obviously follows from the symmetric case. Indeed, for $B<0$, it holds $g(t)\leq g(1-t)$ for $t\in[1/2,1]$, therefore the global minimum lies on the segment $[1/2,1]$. Let s, 1ss,\ 1-s ($0<s<1/2$) be zeros of $g^{\prime\prime}(t)$. On the segment [s,1s][s,1-s] the function $g(t)^{\prime}$ is decreasing, on the segments [0,s],[1s,1][0,s],\ [1-s,1] the function $g(t)^{\prime}$ increases, and due to symmetry $(g(t)+Bt)^{\prime}$ has root $1/2$. Therefore, on the segment $[1/2,1]$ when $B<0$ the function $g(t)$ has no more than one critical point. ∎
Let us finally proceed to the proof of lemma 4. The variables $\alpha,\ p,\ \beta,\ t$ used in what follows, unless otherwise stated, are such that
$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1)},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
Lemma 4.
For $\alpha,\ p,\ \beta,\ t$ specified above one has the inequality
$\mu_{0}(p,\alpha)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\mu_{1}(p,\alpha)$
$\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}.$
Proof.
Let
$f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl[\mu_{0}(p,\alpha)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\mu_{1}(p,\alpha)\Bigl],$
for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) of lemma 3.
The derivative of $f(t)$ is
$f^{\prime}(t)=\log\dfrac{1-t}{t}+\log\dfrac{\beta}{1-\beta}-\mu_{0}(p,\alpha)\Bigg(\sqrt{\dfrac{\beta}{t}}-\sqrt{\dfrac{1-\beta}{1-t}}\Bigg)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl).$
Consider the following cases:
  1. 1.
    $B\leq 0$.
    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to lemma 3 presented later. One has
    $f(\beta)=\mu_{1}(p,\alpha)-\mu_{0}(p,\alpha)\geq 0.$ (90)
    Indeed, the inequality (90) is equivalent to
    $\Bigl(\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}\Bigl)\Bigl(\sqrt{p\alpha}+\sqrt{(1-p)(1-\alpha)}\Bigl)$-
    $-\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)\geq 0\quad\Leftrightarrow$
    $p\log\dfrac{p}{\alpha}\Bigg(1-\sqrt{\dfrac{\alpha(1-p)}{p(1-\alpha)}}\Bigg)-(1-p)\log\dfrac{1-p}{1-\alpha}\Bigg(1-\sqrt{\dfrac{p(1-\alpha)}{\alpha(1-p)}}\Bigg)\geq 0\quad\Leftrightarrow$
    $\sqrt{\alpha p}\log\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\log\dfrac{1-p}{1-\alpha}\geq 0.\quad$ (91)
    Let $V(p)=\sqrt{\alpha p}\ln\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\ln\dfrac{1-p}{1-\alpha}$, then
    1. (a)
      $(d/dp)^{2}V(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
      (since $p\geq 1-\alpha\geq 1/2$);
    2. (b)
      $V(1-\alpha)=0$ and
      $V^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
      $\dfrac{\alpha}{1-\alpha}\Bigl(1+\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl)-\Bigl(1-\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl)\geq 0\quad\Leftrightarrow$
      $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha.$
    Therefore, $V(p)\geq 0$ and the inequality (91) is established.
  2. 2.
    $B\geq 0$.
    This condition implies that the function $f(t)$ has the minimum point $t_{*}$ on $[1/2,1]$ (lemma 3 is applied in the same way). This point is characterized by the equation
    $\log\dfrac{1-t_{*}}{t_{*}}+\log\dfrac{\beta}{1-\beta}-\mu_{0}(p,\alpha)\Bigg(\sqrt{\dfrac{\beta}{t_{*}}}-\sqrt{\dfrac{1-\beta}{1-t_{*}}}\Bigg)\Bigl(\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl)=0.$
    The value of $t_{*}$ depends on $\beta$, and for the function $f(\beta,t)\equiv f(t)$ it holds $\dfrac{d}{d\beta}f(\beta,t_{*}(\beta))=\dfrac{\partial}{\partial\beta}f(\beta,t_{*}(\beta))+\dfrac{dt_{*}(\beta)}{d\beta}\dfrac{\partial}{\partial t}f(\beta,t_{*}(\beta))=\dfrac{\partial}{\partial\beta}f(\beta,t_{*}(\beta))$ since $t_{*}$ is critical for $f(t)$. Then,
    $\dfrac{d}{d\beta}f(t_{*})=\dfrac{1}{\ln 2}\Bigl(\dfrac{t_{*}}{\beta}-\dfrac{1-t_{*}}{1-\beta}\Bigl)-\mu_{0}(p,\alpha)\Bigg(\sqrt{\dfrac{t_{*}}{\beta}}-\sqrt{\dfrac{1-t_{*}}{1-\beta}}\Bigg)\times$
    $\times\Bigl(\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl).$
    Hence
    $\ln 2\ \dfrac{d}{d\beta}f(t_{*})=\Bigl(\dfrac{t_{*}}{\beta}-\dfrac{1-t_{*}}{1-\beta}\Bigl)+\sqrt{\dfrac{t_{*}(1-t_{*})}{\beta(1-\beta)}}\Bigl(\ln\dfrac{1-t_{*}}{t_{*}}-\ln\dfrac{1-\beta}{\beta}\Bigl)\geq 0$
    since
    $\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}-\sqrt{\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}}+\ln\Bigl(\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}\Bigl)=R-1/R-2\ln R\geq 0$
    where $R\equiv\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}\geq 1$.
    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$. We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and hence for all the other points $t\in[0,1]$.
In the special case of (91), where $\alpha=1/2$, one has an interesting inequality
$h_{1/2}(p)\equiv\sqrt{2p}\log(2p)+\sqrt{2(1-p)}\log(2(1-p))\geq 0,$
which takes place for $p\in(0,1)$. The function $h_{1/2}(p)$ is nonnegative, convex and $h_{1/2}(p)\leq\sqrt{2}$.
5.2 The case of theorem 3
This section is devoted to proving the inequality (72) which verifies the condition (ii.a) in the case of obtuse pyramid. Namely, if the parameter $p<1$ satisfies $p\geq\max\left\{\dfrac{m-1}{m},p(m)\right\}$ with $p(m)$ defined in (78), then for complex numbers $z_{1},...,z_{m}$ satisfying (18) the inequality
$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl|\sum\limits_{j=1}^{m}z_{j}\Bigl|^{2}+\tilde{\mu}_{1}(p)\geq 0.$ (92)
holds with $\tilde{\mu}_{0}(p),\tilde{\mu}_{1}(p)$ defined in (73), (74).
Suggestion for proof.
As it has been established in lemma 1, a real case is enough for analysis.
For real vectors $(z_{j})_{j=1}^{m}$ from the unit sphere let us consider the Lagrange function
$L(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}z_{k}^{2}\log z_{k}^{2}-\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{k=1}^{m}z_{k}\Bigl)^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl(\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl).$
The necessary condition is
$\dfrac{1}{2}\dfrac{\partial L}{\partial z_{k}}=z_{k}\log z_{k}+\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}z_{j}\Bigl)+(\lambda+\log e)z_{k}=0,\quad 1\leq k\leq m,$
and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since the equation
$z\log z^{2}+\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}z_{j}\Bigl)+(\lambda+\log e)z=0$ (93)
for a variable $z$ has at most three real solutions. The fact the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then three real roots was actually used in proof of lemma 2 for equation (86) which is similar to (93) with $\leq 2$ roots. However, here we consider the entire real line, so the maximum number of real roots is three.
Consider the following cases.
Case 1. The minimizer has a zero component.
Then $\lambda+\log e=0$ and the equation (93) has roots 0 and $\pm s$. Therefore the minimizer of the function $\tilde{F}(z)$ should have the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$. The value $1+\tilde{\mu}_{1}(p)$ is nonnegative for the following reason.
First we will prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl[\dfrac{m-1}{m},1\Bigr)$. The derivative is
$\tilde{\mu}_{1}^{\prime}(p)=\log e\dfrac{2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}-4p+2+\ln p-\ln\dfrac{1-p}{m-1}}{2(m-1)\sqrt{\dfrac{p(1-p)}{m-1}}\Bigg(\sqrt{p}+\sqrt{\dfrac{1-p}{m-1}}\Bigg)^{2}}.$
Here $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, therefore, $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
Thus for $\frac{m-1}{m}\leq p<1$ we have
$\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}\Bigl(\frac{m-1}{m}\Bigl)=-\dfrac{m-2}{m}\log(m-1)+\log m$
due to (75). For $p(m)\leq p<1$ we have
$\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
because of (79). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
Case 2. The minimizer has no zero components.
To reduce the number of different values to two, we will introduce a new Lagrange function $\hat{L}$. The critical points of $L$ are related with critical points of the Lagrange function
$\hat{L}(s_{1},s_{2},s_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}\log s_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}k_{j}s_{j}\Bigl)^{2}$-
$-\lambda\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}-1\Bigl)-\nu\Bigl(\sum\limits_{j=1}^{3}k_{j}-m\Bigl).$
New necessary conditions are
$-\dfrac{1}{2}\dfrac{\partial\hat{L}}{\partial s_{l}}=k_{l}s_{l}\log s_{l}^{2}+\tilde{\mu}_{0}(p)k_{l}\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+(\lambda+\log e)k_{l}s_{l}=0,$
$-\dfrac{\partial\hat{L}}{\partial k_{l}}=s_{l}^{2}\log s_{l}^{2}+2\tilde{\mu}_{0}(p)s_{l}\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+\lambda s_{l}^{2}+\nu=0$
for all $1\leq l\leq 3$.
Thus, we have
$\dfrac{1}{2}s_{l}\dfrac{\partial\hat{L}}{\partial s_{l}}+k_{l}\dfrac{\partial\hat{L}}{\partial k_{l}}=k_{l}s_{l}\Bigl(s_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)-\nu k_{l}=0.$
Case 2.1. $k_{1},k_{2},k_{3}\neq 0$.
For all $1\leq l\leq 3$ it holds $s_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+\nu$, which can take place, if the set $\{s_{1},s_{2},s_{3}\}$ has no more than two different values (which is realized, formally, for fractional multiplicities). This can be regarded as the next case.
Case 2.2. $k_{3}=0$.
Up to permutation and the global sign, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}>0$}, $s_{2}<0$}. The components of the minimizer should have different signs; otherwise, changing the sign of a single component would yield a lower value for $\tilde{F}$.
Denote
$\tilde{\mu}_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)}{\Bigl(\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}+\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$
$\tilde{\mu}_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}+\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}+\sqrt{\dfrac{1-p}{1-\alpha}}},$
where $\alpha$ and $p$ lie on $(0,1)$.
Relation with $\tilde{\mu}_{0}(p)$ and $\tilde{\mu}_{1}(p)$ is given by
$\tilde{\mu}_{0}(p)=\dfrac{1}{m}\tilde{\mu}_{0}(p,1/m),$
$\tilde{\mu}_{1}(p)=\tilde{\mu}_{1}(p,1/m)-\log m.$
It is easy to see that for $\alpha$ and $p$ such that $0<\alpha\leq 1/2\leq 1-\alpha\leq p<1$ one has $\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\geq 0$ and $\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\leq 0$}, therefore, $\tilde{\mu}_{0}(p,\alpha)\leq 0$.
The initial inequality (72) for $k_{1}s_{1}^{2}=t$ is equivalent to
$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl(\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl)^{2}-\tilde{\mu}_{1}(p,1/m),$ (94)
and the statement follows from lemma 5.
Lemma 5.
Let $0\leq\alpha\leq\beta\leq\dfrac{1}{2}\leq 1-\alpha\leq p\leq 1$, $0\leq t\leq 1$. Then
$\tilde{f}(t)\equiv-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigg(\tilde{\mu}_{0}(p,\alpha)\Bigl(\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\tilde{\mu}_{1}(p,\alpha)\Bigg)\geq 0.$
Proof.
We will prove that the inequality
$\mu_{0}(p,\alpha)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\mu_{1}(p,\alpha)\geq\tilde{\mu}_{0}(p,\alpha)\Bigl(\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\tilde{\mu}_{1}(p,\alpha)$
holds in this range. So, the statement of the lemma follows from lemma 4.
Collecting similar terms and using notation of $f(t)$ from subsection 5.1,
$\tilde{f}(t)-f(t)=2\sqrt{\beta(1-\beta)}\Bigl(\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl)\sqrt{t(1-t)}-(1-2\beta)\Bigl(\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl)t+$
$+\Bigl[(1-\beta)\Bigl(\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl)-\Bigl(\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)\Bigl)\Bigl]\geq 0.$
One can see that
$\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)=2(2p-1)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)}{(p-\alpha)(p+\alpha-1)},$
$\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)=-2(1-2\alpha)p(1-p)\dfrac{\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}}{(p-\alpha)(p+\alpha-1)}\leq 0,$
$\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)=2\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)}{p-\alpha}.$
The function $(\tilde{f}-f)(t)$ has the form $A\sqrt{t(1-t)}+Bt+C$ with $A\leq 0$ and $B\leq 0$, then the unique minimum point $t_{*}\in[1/2,1]$. It has the value $t_{*}=\dfrac{1}{2}\Bigl(1-\dfrac{B}{\sqrt{A^{2}+B^{2}}}\Bigl)$ and the minimum of the function $(\tilde{f}-f)(t)$ is equal to
$(\tilde{f}-f)(t_{*})=\dfrac{B}{2}+C-\dfrac{1}{2}\sqrt{A^{2}+B^{2}},$
where
$A=2\sqrt{\beta(1-\beta)}\Bigl(\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl),$
$B=-(1-2\beta)\Bigl(\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl),$
$C=\Bigl[(1-\beta)\Bigl(\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl)-\Bigl(\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)\Bigl)\Bigl].$
One has
$\dfrac{B}{2}+C=\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)}{(p-\alpha)(p+\alpha-1)}\times$
$\times\Bigl((2p-1)-2(p+\alpha-1)\Bigl)=$
$=(1-2\alpha)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)}{(p-\alpha)(p+\alpha-1)}\geq 0,$
$\dfrac{1}{2}\sqrt{A^{2}+B^{2}}=(1-2\alpha)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl)}{(p-\alpha)(p+\alpha-1)}\times$
$\times\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}$
Thus,
$\dfrac{\sqrt{A^{2}+B^{2}}}{B+2C}=\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}\leq 1$ (95)
Let us denote u(t)=4t(1t)u(t)=4t(1-t). The inequality (95) can be rewritten in the form
$\dfrac{u(\beta)u(p)}{u(\alpha)}+\dfrac{(1-u(\beta))(1-u(p))}{1-u(\alpha)}\leq 1,$
where $u(p)\leq u(\alpha)\leq u(\beta)$. The linear function
$t\mapsto W(t)=\dfrac{u(\beta)t}{u(\alpha)}+\dfrac{(1-u(\beta))(1-t)}{1-u(\alpha)}$
has non-negative derivative $\dfrac{u(\beta)}{u(\alpha)}-\dfrac{1-u(\beta)}{1-u(\alpha)}=\dfrac{u(\beta)-u(\alpha)}{u(\alpha)(1-u(\alpha))}$ and $W(u(\alpha))=1$}. Therefore, the inequality (95) holds, and lemma has been proven. ∎
Acknowledgment. The authors are grateful to S.V. Grishin and D.A. Kronberg for providing numerical simulations.